# MIPHEI-vit: Multiplex Immunofluorescence Prediction from H&E Images using ViT Foundation Models

### v1

@@ 7 Discussion
In this study, we present MIPHEI, a method trained to predict 16 mIF channels from standard H&E slides. For this, we proposed a U-Net with a ViT-based foundation model as encoder and demonstrated that this architecture outperforms previously proposed methods. Moreover, we showed that MIPHEI generalizes well across datasets. We attribute this to the robust and transferable encodings learned by the foundation model, which was exposed to millions of tiles during pretraining. Our findings highlight the value of leveraging foundation models for mIF prediction, and potentially for other image translation tasks in histology.
We also introduced a validation strategy focused on single-cell metrics, which we believe are the most relevant for this task. Pixel-wise accuracy may be unreliable, as H&E images are usually not informative about cytoplasmic boundaries. Instead, the biologically meaningful information lies in protein expression levels within individual cells, protein positivity, or the resulting cell type. We reflect this in our evaluation framework, which we provide as part of this study.
Accurate, domain-robust mIF prediction opens the door to a range of applications. While direct clinical deployment for diagnostics may remain challenging, MIPHEI proves to be a powerful tool for mining large retrospective cohorts—enabling the identification of cell types without relying on manual annotation. Ultimately, this can allow to identify associations between specific cell populations, their spatial arrangements, and clinically relevant outcomes such as survival or treatment response. MIPHEI thus holds promise for hypothesis generation and exploratory analysis. Furthermore, extending this approach to predict clinically relevant scores, such as the Immunoscore, represents a compelling direction for future work.
@@ 8 Conclusion
In this paper, we present MIPHEI, a deep learning framework that predicts mIF images from H&E using ViT foundation models as encoders within a U-Net architecture. MIPHEI outperforms state-of-the-art models on both internal and external datasets, demonstrating strong generalization across staining protocols and imaging conditions. We evaluated MIPHEI across 15 protein markers and associated cell types. It achieves high accuracy for epithelial (Pan-CK, E-cadherin) and broad immune markers (CD45, CD3e), performs moderately on more specific immune subtypes (CD8a, CD4, CD45RO, CD68, CD163) and stromal markers (CD31, $\alpha$-SMA), and struggles with FOXP3 and PD-L1 due to their complexity and small number of positive cells.
@@ Acknowledgment
Funding
This work was sponsored by ANRT and Sanofi. Furthermore, T. Walter acknowledges funding from Agence Nationale de la Recherche under the France 2030 program, with the reference number ANR-24-EXCI-0004
Conflict of Interest

### v2

@@ 7 Discussion
In this study, we present MIPHEI, a method trained to predict 16 mIF channels from standard H&E slides. For this, we proposed a U-Net with a ViT-based foundation model as encoder and demonstrated that this architecture outperforms previously proposed methods. Moreover, we showed that MIPHEI generalizes well across datasets. We attribute this to the robust and transferable encodings learned by the foundation model, which was exposed to millions of tiles during pretraining. This contrasts with prior reports (Andani et al., 2025; Wu et al., 2025), where foundation models underperformed, likely due to limited fine-tuning strategies or insufficient integration within the architecture. Our findings highlight the value of leveraging foundation models for mIF prediction, and potentially for other image translation tasks in histology.
We also introduced a validation strategy focused on single-cell metrics, which we believe are the most relevant for this task. Pixel-wise accuracy may be unreliable, as H&E images are usually not informative about cytoplasmic boundaries, and mIF targets are often sparse. Instead, the biologically meaningful information lies in protein expression levels within individual cells, protein positivity, or the resulting cell type. We reflect this in our evaluation framework, which we provide as part of this study.
Accurate, domain-robust mIF prediction opens the door to a range of applications. While direct clinical deployment for diagnostics may remain challenging, MIPHEI proves to be a powerful tool for mining large retrospective cohorts, enabling the identification of cell types without relying on manual annotation. Ultimately, this can allow to identify associations between specific cell populations, their spatial arrangements, and clinically relevant outcomes such as survival or treatment response. MIPHEI thus holds promise for hypothesis generation and exploratory analysis. Furthermore, extending this approach to predict clinically relevant scores, such as the Immunoscore, represents a compelling direction for future work.
@@ 8 Conclusion
In this paper, we present MIPHEI, a deep learning framework that predicts mIF images from H&E using ViT foundation models as encoders within a U-Net architecture. MIPHEI outperforms state-of-the-art models on both internal and external datasets, demonstrating strong generalization across staining protocols and imaging conditions. We evaluated MIPHEI across 15 protein markers and associated cell types. On our in-domain test set, it achieves high accuracy for epithelial (Pan-CK, E-cadherin), broad immune markers (CD45, CD3e) and stromal markers ($\alpha$-SMA), performs moderately on more specific immune subtypes (CD4, CD45RO, CD20, CD68) and vascular markers (CD31), and struggles with CD8a, FOXP3 and PD-L1 due to their complexity and small number of positive cells.
CRediT authorship contribution statement
Guillaume Balezo: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Validation, Visualization, Writing – original draft, Writing – review & editing. Roger Trullo: Conceptualization, Funding acquisition, Methodology, Project administration, Resources, Supervision, Writing – review & editing. Albert Pla Planas: Conceptualization, Methodology, Project administration, Resources, Supervision, Visualization, Writing – review & editing. Etienne Decenciere: Conceptualization, Methodology, Project administration, Supervision, Visualization, Writing – review & editing. Thomas Walter: Conceptualization, Funding acquisition, Methodology, Project administration, Supervision, Visualization, Writing – original draft, Writing – review & editing.
Declaration of competing interest
@@ Acknowledgment
This work was sponsored by ANRT and Sanofi. Furthermore, T. Walter acknowledges funding from Agence Nationale de la Recherche under the France 2030 program, with the reference numbers ANR-24-EXCI-0001, ANR-24-EXCI-0002, ANR-24-EXCI-0003, ANR-24-EXCI-0004, ANR-24-EXCI-0005, as well as funding by the French government under the management of Agence Nationale de la Recherche as part of the “Investissements d’avenir” program, reference ANR-19-P3IA-0001 (PRAIRIE 3IA Institute) and ANR-23-IACL-0008 (PRAIRIE-PSAI).
Data availability
The dataset used in this study, derived from the OrionCRC and HEMIT datasets, has been made publicly available on Zenodo at https://doi.org/10.5281/zenodo.15340874 (Balezo et al., 2025). Only the IMMUcan data used for validation is not redistributed, as it remains private and subject to usage restrictions.
Declaration of generative AI and AI-assisted technologies in the writing process
We used ChatGPT (OpenAI) exclusively to improve the English language quality of the manuscript—specifically for grammar, spelling, and phrasing. It was also used to assist with formatting and documentation of the experimental code. All content was written by the authors, and the final text was carefully reviewed to ensure accuracy and clarity.